# Federated Learning Privacy: Attacks, Defenses, Applications, and Policy Landscape — A Survey

### v1

@@ Abstract.
Deep learning has shown incredible potential across a vast array of tasks and accompanying this growth has been an insatiable appetite for data. However, a large amount of data needed for enabling deep learning is stored on personal devices and recent concerns on privacy have further highlighted challenges for accessing such data. As a result, federated learning (FL) has emerged as an important privacy-preserving technology enabling collaborative training of machine learning models without the need to send the raw, potentially sensitive, data to a central server. However, the fundamental premise that sending model updates to a server is privacy-preserving only holds if the updates cannot be ”reverse engineered” to infer information about the private training data. It has been shown under a wide variety of settings that this premise for privacy does not hold.
In this survey paper, we provide a comprehensive literature review of the different privacy attacks and defense methods in FL. We identify the current limitations of these attacks and highlight the settings in which FL client privacy can be broken. We dissect some of the successful industry applications of FL and draw lessons for future successful adoption. We survey the emerging landscape of privacy regulation for FL. We conclude with future directions for taking FL toward the cherished goal of generating accurate models while preserving the privacy of the data from its participants.
@@ 1. Introduction
Federated Learning (FL) is a widely popular learning architecture that allows one to learn a Machine Learning (ML) model collaboratively. The classical structure of FL is that there are multiple clients each with their own local data, which they would possibly like to keep private, and there is a server that is responsible for learning a global ML model.
One of the two primary reasons for the popularity of FL is that clients can keep their data private and still benefit from the combined learning across all of their data. (The second reason is the “power of crowds”, i.e., many weak devices can come together to learn complex models, which would be beyond the compute power of any one client to learn on its own.) However, by 2018, this hope of FL had been effectively dashed. It was shown in a set of seminal papers that if the central aggregator has access to gradients sent by the clients (which it does in most versions of FL), then the aggregator can learn various things from these clients, each of which would be taken to effectively break the privacy of client data. (Zhu et al., 2019; Zhao et al., 2020; Melis et al., 2019) The simplest to understand form of this attack is that the aggregator can reconstruct the data of the clients from the gradient updates, to different degrees of fidelity. The attacker, the central aggregator here, has the unique advantage that it can observe the individual updates from the clients over time and can control the view of the participants of the global parameters. Also worryingly, the attacker could be one of the clients who can observe the global parameter updates, and can control his parameter uploads.
In this article, we survey the constant back and forth that has been going on for the last 5 years or so in protecting the privacy of data through FL. We will have as one of the primary focus questions: what are the research challenges that need to be solved to achieve “good enough” protection for different application use cases. Our article takes a distinctive viewpoint of grounding the technical discussion of the attacks and the defenses in application use cases. For example, the kind of privacy expectation, and regulation, in the healthcare sector is very different from that in the finance or insurance sector, which again is distinct from that in Internet of Things (IoT).
@@
Policy considerations. Our article goes beyond the technology landscape and also surveys the policy landscape (Section 5), which is nascent and developing. A central consideration of ML regulation, namely, regulation of the data, is central to FL since here we deal with data that is local, and possibly sensitive, to individual clients. Regulation of FL is therefore apt and a start needs to be made soon. We hope that this discussion of regulation, in the context of the technical achievements to date and the expected deliverables, will lead us to enlightened regulation in the space of privacy in FL.
Privacy pushes-and-pulls. Given that being a privacy-preserving technology was a core reason for its introduction, it is vital that FL is indeed able to preserve the privacy of users. Despite this, the process of sending gradient or model updates to a server has been shown to be less than secure. Recent works in membership inference, property inference, and even data reconstruction have demonstrated that vanilla FL is vulnerable to a multitude of privacy attacks. On the other hand, defenses such as secure aggregation, differential privacy, or homomorphic encryption have also continued to be developed to prevent these attacks. Ultimately, the privacy expectation of FL is that an attacker, whether being the server, an outsider, or a participating client, is not able to infer any private information about the local training data of any of the clients. In this work, we discuss the current privacy attacks and pinpoint the vulnerabilities that make them possible. We then highlight the defenses that can prevent the attacks or make them more difficult along with the limitations and drawbacks of each.
@@ 2.1. Federated learning process
FL has emerged as a decentralized training framework that addresses the dual challenges of leveraging computational resources at the network’s edge and preserving data privacy. Central to FL is its unique approach to collaborative model training, which aggregates model updates from participating clients rather than transmitting the raw data of the clients. This method not only leverages the vast computational power distributed across edge devices but also aligns with the increasing demands for data privacy amidst the exponential growth and insatiable appetite for data of modern AI systems. At the heart of FL’s philosophy is the distributed training of machine learning models, where the crux is the exchange of model updates with a central server, sidestepping the need to share sensitive training data.
The vanilla FL training process typically involves the coordinated interplay of two principal actors: the central server and the participating clients. The central server orchestrates the workflow by selecting the clients, dispatching the global model to the participating clients for local refinement, and then aggregating the updates before finally updating the global model. Conversely, clients engage in a straightforward yet crucial role. They receive the global model from the server, apply training updates using their local datasets, and then transmit these training updates back to the server.
For a single FL training cycle, the operations unfold as follows:
@@
In cross-device FL, typically involving lightweight devices like smartphones or IoT devices, the selection ranges from to . This broad participation mitigates network unreliability and device dropouts, ensuring diverse data representation.
@@
Aggregation and Global Model Update: The server aggregates these client updates using methods like mean or weighted mean to form the new global model parameters, . This prepares the model for the next distribution, refinement, and aggregation cycle.
@@
On the receiving end of this process, the central server plays a critical role in aggregating the updates from all participating clients. Given the potential for compromised devices or malicious actors within the network, the server might implement robust aggregation techniques. These techniques are designed to identify and neutralize the influence of any anomalous updates that could skew the model’s learning in an undesirable direction. By employing such strategies, the server ensures that the aggregated update accurately reflects the collective learning from all legitimate participants, thereby reinforcing the reliability and effectiveness of the FL process.
These considerations highlight the need for a multi-layered approach to address the practical challenges in deploying secure and private FL systems. There is often a tension between operational efficiency and the security/privacy achieved by the protocol.
@@ 2.3.1. Cross-device and Cross-silo Federated Learning:
FL can be broadly categorized into cross-device and cross-silo contexts, each presenting unique scalability and deployment challenges. In cross-device FL, we envisage a vast network of personal devices, such as smartphones, wearables, or IoT devices, each participating in the learning process. This variant is marked by its massive scale, involving potentially millions of devices, each contributing data for model training. The finance sector, for example, could leverage cross-device FL to enhance fraud detection systems by learning from transactions across countless mobile banking apps without centralizing sensitive financial data. However, this approach is fraught with challenges, primarily due to the limited computational power and intermittent connectivity of devices, not to mention the heightened concerns over data privacy and security. Conversely, cross-silo FL operates on a smaller scale with fewer, resource-ample entities like hospitals, featuring stable communications and higher computational resources. n healthcare, for instance, cross-silo FL enables hospitals to collaboratively improve diagnostic models by learning from diverse patient datasets while ensuring that sensitive medical records remain within hospital premises. Similarly, in finance, banks can utilize cross-silo FL to jointly develop more accurate credit scoring models without exposing their clients’ financial details to competitors. This form of FL, while offering more control and resources, still necessitates careful navigation of regulatory, privacy, and interoperability challenges. FL showcases its adaptability and value across sectors through its two main variants: cross-device and cross-silo. In finance, cross-device FL can harness personal devices like smartphones to enhance fraud detection without centralizing sensitive data, balancing privacy with regulatory adherence despite challenges like limited device power and connectivity. Meanwhile, healthcare can benefit from cross-silo FL, where institutions like hospitals collaborate to improve diagnostics without sharing patient data, aligning with strict privacy laws and ethical standards. These approaches demonstrate FL’s capacity to meet sector-specific needs—offering scalable, privacy-preserving solutions in finance through cross-device models, and secure, collaborative advancements in healthcare via cross-silo frameworks. This duality not only underscores FL’s versatility in tackling regulatory and privacy challenges but also emphasizes its potential to drive innovation and security in sensitive domains.
@@ 2.3.2. Federated transfer learning:
The crux of FTL’s effectiveness lies in its domain adaptation strategies, where the challenge is to make a model trained in one domain perform well in another with distinct data distributions, feature spaces, and tasks. This process, inherently more complex than the simple update aggregation in FL, demands sophisticated communication for exchanging not just model updates but also domain-specific adaptations, leading to higher communication overhead. Moreover, FTL must navigate the delicate balance between privacy and utility. It leverages the privacy-preserving nature of FL by ensuring that sensitive data remains within its original domain, with additional measures like differential privacy further enhancing security. However, this focus on privacy necessitates careful handling to ensure that the transferred model knowledge does not inadvertently leak sensitive information.
FTL’s requirement for domain adaptation introduces significant communication overhead and computational complexity. The iterative process of model adaptation and fine-tuning across domains entails multiple rounds of communication, each adding to the logistical and computational demands of deploying FTL at scale. Despite these challenges, the potential of FTL to leverage cross-domain knowledge without compromising data privacy stands as a compelling advancement in distributed machine learning.
FTL is particularly useful in scenarios where data cannot be pooled together due to privacy concerns but where tasks across domains are related. For example, in healthcare, models trained on data from one hospital can be adapted to improve predictions in another hospital, despite differences in patient demographics or data collection practices. Similarly, in finance, knowledge from credit scoring models in one region can be transferred to enhance models in another region.
@@ 2.3.3. Hierarchical Federated Learning
Hierarchical Federated Learning (HFL) (Abad et al., 2020; Feng et al., 2021; Liu et al., 2020; Briggs et al., 2020; Liu et al., 2022b) introduces a multi-tiered aggregation framework, significantly enhancing the scalability and efficiency of collaborative learning models. This system incorporates several levels of servers, from edge parameter servers to a final cloud parameter server, facilitating a more structured and efficient communication and aggregation pathway. By leveraging edge servers for initial aggregation (Liu et al., 2020), HFL mitigates the communication bottlenecks often encountered in traditional FL, enabling faster, more energy-efficient training processes. In response to the limitations observed in both FL and P2PL — where FL’s centralized model can pose privacy risks, and P2PL’s decentralized nature complicates integrity and consensus — HFL emerges as a powerful hybrid solution. It employs a hierarchical structure where clients, or “spokes,” communicate updates to edge servers, or “hubs,” which then aggregate these updates before further communication among themselves or with a higher-level server. This architecture not only reduces the privacy concerns associated with a single central server but also addresses the communication and consensus challenges inherent in a fully decentralized setup.
Horizontal and Vertical Federated Learning
@@
2.3.4. IID and non-IID Data Distributions
IID Data Distributions: In scenarios where data is considered IID, each client’s dataset is assumed to be drawn from the same probability distribution. This uniformity suggests that datasets across varied domains, such as those of a bank and a hospital, exhibit similar statistical properties and structures, despite their distinct contexts.
Non-IID Data Distributions: Contrastingly, non-IID data distributions reflect situations where data across clients originate from heterogeneous distributions. This diversity is rooted in demographic variances, regional characteristics, or the intrinsic nature of the collected data.
Implications on Privacy and Learning: Privacy Concerns: The heterogeneity of non-IID data amplifies privacy concerns, necessitating robust privacy-preserving mechanisms that can handle the diversity of data sources. The specificity of models to certain data characteristics could inadvertently lead to data leakage or facilitate inference attacks, underscoring the need for privacy techniques that are adaptable to varying data distributions.
@@ 2.3.5. FedSGD and FedAvg: Core FL algorithms
The two standard FL algorithms are FedSGD and FedAvg (McMahan et al., 2017), mainly differing in the number of local iterations. FedSGD, the federated adaptation of stochastic gradient descent, operates on the principle of minimalism. It mandates that clients engage in a singular iteration of training on a localized data batch before transmitting the resultant gradient to a central server. This server, in turn, aggregates the gradients from all participating clients, adjusting the weights in proportion to each client’s contribution to the total dataset. This process is especially pertinent in scenarios demanding rapid model updates.
Advancing beyond FedSGD, FedAvg introduces multiple local iterations or epochs at each client before communicating updates to the server. These updates can comprise either the computed gradients over all local iterations or the updated model parameters themselves. The algorithm’s structure inherently minimizes the frequency of communication with the server, addressing privacy concerns by limiting the exposure of initial and final model states only. This characteristic is crucial in contexts where data sensitivity is paramount, as it effectively conceals the intermediate computational states from potential external eavesdropping. Moreover, FedAvg facilitates a deeper, more nuanced learning from the complex datasets typical of financial and medical sectors. By allowing for extended local training periods, the algorithm empowers the model to better capture and adapt to the intricate patterns present within such diverse data sources. This capability not only enhances the model’s precision but also its applicability and relevance to the distinct challenges and requirements of these domains.
@@ 2.4. Threat model
For privacy attacks within the domain of FL, the the threat landscape is complex and varies significantly based on the nature of the attack. The architecture of FL inherently poses unique challenges in safeguarding privacy due to the decentralized nature of data processing. Here, the central server and the clients play pivotal roles, each representing potential vectors for privacy breaches and attacks. Understanding the nuances of these threats is crucial for developing robust defense mechanisms. At the heart of FL, the central server embodies a critical point of vulnerability. It plays a dual role: distributing the learning model to clients and aggregating their updates. This centralized control positions the server as a potent target for attacks, making it the focal point for the strongest privacy threats in FL. The server’s pivotal role in managing the flow of information — both disseminating and receiving — means that attacks can take on various forms, each tailored to exploit specific weaknesses in the server-client interaction.
@@
Complexities of Misbehaving Clients: The threat landscape is further complicated by the variability in the behavior of misbehaving clients. Two critical factors need consideration: the proportion of clients engaged in malicious activities and the extent of their collaboration. Misbehaving clients may operate independently, presenting a scattered threat. However, the danger amplifies when they collude, orchestrating their attacks to maximize disruption or data extraction. This coordinated effort can significantly exacerbate the challenges in defending against privacy breaches and ensuring the integrity of the FL process.

### v2

# The Federation Strikes Back: A Survey of Federated Learning Privacy Attacks, Defenses, Applications, and Policy Landscape

@@ Abstract.
Deep learning has shown incredible potential across a wide array of tasks, and accompanied by this growth has been an insatiable appetite for data. However, a large amount of data needed for enabling deep learning is stored on personal devices, and recent concerns on privacy have further highlighted challenges for accessing such data. As a result, federated learning (FL) has emerged as an important privacy-preserving technology that enables collaborative training of machine learning models without the need to send the raw, potentially sensitive, data to a central server. However, the fundamental premise that sending model updates to a server is privacy-preserving only holds if the updates cannot be ”reverse engineered” to infer information about the private training data. It has been shown under a wide variety of settings that this privacy premise does not hold.
In this survey paper, we provide a comprehensive literature review of the different privacy attacks and defense methods in FL. We identify the current limitations of these attacks and highlight the settings in which the privacy of ann FL client can be broken. We further dissect some of the successful industry applications of FL and draw lessons for future successful adoption. We survey the emerging landscape of privacy regulation for FL and conclude with future directions for taking FL toward the cherished goal of generating accurate models while preserving the privacy of the data from its participants.
@@ 1. Introduction
Federated Learning (FL) is a popular learning paradigm that allows one to learn a Machine Learning (ML) model collaboratively. The classical structure of FL is with multiple clients each having their own local data, which they would possibly like to keep private, and there is a server that is responsible for learning a global ML model. 111The phrase “The Federation Strikes Back” is a playful nod to the The Empire Strikes Back, symbolizing the ongoing arms race between privacy attackers and defenders in the federated learning domain, where data confidentiality stands at the heart of the conflict.
One of the two primary reasons for the popularity of FL is that clients can keep their data private and still benefit from the combined learning across all of their data. (A second reason is the “power of crowds”, i.e., many weak devices can come together to learn complex models, which would be beyond the compute power of any one client to learn on its own.) However, by 2018, the privacy of FL has been put into question questioned. In a set of seminal papers, it was shown that a central aggregator with access to gradients sent by the clients (which is standard in most versions of FL), then the aggregator can learn various things from these clients, each of which would be taken to effectively break the privacy of client data. (Zhu et al., 2019; Zhao et al., 2020; Melis et al., 2019) The simplest form of this attack is where the aggregator can reconstruct the data of the clients from the gradient updates, to different degrees of fidelity. The attacker, the central aggregator here, has the unique advantage that it can observe individual updates from the clients over time and can control the view of the participants of the global parameters. Also worryingly, the attacker can even be one of the clients who can observe the global parameter updates, and can control his parameter uploads (Hitaj et al., 2017).
In this article, we survey the constant back and forth that has been roughly going on for the last 5 years in protecting the privacy of data through FL. We will have as one of the primary focus questions: what are the research challenges that need to be solved to achieve “good enough” protection for different application use cases. Our article takes a distinctive viewpoint of grounding the technical discussion of the attacks and the defenses in application use cases. For example, the kind of privacy expectation, and regulation, in the healthcare sector is very different from that in the finance or insurance sector, which again is distinct from that in Internet of Things (IoT).
@@
Policy considerations. Our article goes beyond the technology landscape and also surveys the policy landscape (Section 5), which is still developing. A central consideration of ML regulation, namely, regulation of the data, is central to FL since here we deal with data that is local, and possibly sensitive, to individual clients. Regulation of FL is therefore important and a start needs to be made soon. We hope that this discussion of regulation, in the context of the technical achievements to date and the expected deliverables, will lead to better regulation in the space of privacy in FL.
Privacy pushes-and-pulls. Given that being a privacy-preserving technology was a core reason for its introduction, it is vital that FL is indeed able to preserve the privacy of users. However, the process of sending gradients or model updates to a server has been shown to be less than secure. Recent works in membership inference, property inference, and even data reconstruction have demonstrated that vanilla FL is vulnerable to a multitude of privacy attacks. Defenses such as secure aggregation, differential privacy, or homomorphic encryption have also continued to be developed to prevent these attacks. Ultimately, the privacy expectation of FL is that an attacker, whether it is the server, an outsider, or a participating client, is not able to infer any private information about the local training data of any of the clients. In this work, we discuss current privacy attacks and pinpoint the vulnerabilities that make them possible. We then highlight the defenses that can prevent the attacks or make them more difficult along with the limitations and drawbacks of each.
@@ 2.1. Federated learning process
FL has emerged as a decentralized training framework that addresses the dual challenges of leveraging computational resources at the network’s edge and preserving data privacy. Central to FL is its unique approach to collaborative model training, which aggregates model updates from participating clients rather than transmitting the raw data of the clients. This method not only leverages the vast computational power distributed across edge devices but also aligns with the increasing demands for data privacy amidst the exponential growth and insatiable appetite for data of modern AI systems. At the heart of FL is the distributed training of machine learning models, where the crux lies in the exchange of model updates with a central server, sidestepping the need to share sensitive training data.
The vanilla FL training process typically involves the coordinated interplay of two principal actors: the central server and the participating clients. The central server orchestrates the workflow by selecting the clients, dispatching the global model to the participating clients for local refinement, and then aggregating the updates before finally updating the global model again. Conversely, clients engage in a straightforward yet crucial role. They receive the global model from the server, apply training updates using their local datasets, and then transmit their updates back to the server.
For an FL training cycle, the operations unfold as follows:
@@
In cross-device FL, typically involving lightweight devices like smartphones or IoT devices, the selection ranges from s to s. Device communication and computation are typically more unreliable.
@@
Aggregation and Global Model Update: The server aggregates these client updates using methods such as a weighted mean or other aggregation method to form the new global model parameters, . This prepares the model for the next distribution, refinement, and aggregation cycle.
@@
The central server plays a critical role in aggregating the updates from all participating clients. Given the potential for compromised devices or malicious actors within the network, the server might implement robust aggregation techniques (Sharma et al., 2021, 2023a). These techniques are designed to identify and neutralize the influence of any anomalous updates that could skew the model’s learning in an undesirable direction. By employing such strategies, the server ensures that the aggregated update accurately reflects the collective learning from all legitimate participants, thereby reinforcing the reliability and effectiveness of the process.
@@ 2.3.1. Cross-device and Cross-silo Federated Learning:
FL can be broadly categorized into cross-device and cross-silo contexts, each presenting unique scalability and deployment challenges. Cross-device FL typically operates with a vast network of personal devices, such as smartphones, wearables, or IoT devices, each participating in the learning process. This variant is marked by its massive scale, involving potentially millions of devices, each contributing data for model training. The finance sector, for example, could leverage FL to enhance fraud detection systems by learning from transactions across countless mobile banking apps without centralizing sensitive financial data. However, cross-device FL still has many challenges, primarily due to the limited computational power and intermittent connectivity of devices, not to mention the heightened concerns over data privacy and security.
Conversely, cross-silo FL operates on a smaller scale with fewer, resource-ample entities like hospitals, featuring stable communications and higher computational resources. In healthcare, for instance, cross-silo FL enables hospitals to collaboratively improve diagnostic models by learning from diverse patient datasets while ensuring that sensitive medical records remain within hospital premises. Similarly, in finance, banks can utilize cross-silo FL to jointly develop more accurate credit scoring models without exposing their clients’ financial details to competitors. This form of FL, while offering more control and resources, still necessitates careful navigation of regulatory, privacy, and interoperability challenges. We discuss the applications of various privacy-preserving defenses for each category in-depth in Section 7.
@@ 2.3.2. Federated transfer learning:
FTL is particularly useful in scenarios where data cannot be pooled together due to privacy concerns but where tasks across domains are related. For example, in healthcare, models trained on patient data from one hospital on one task can be adapted to improve predictions in another hospital on a different task, despite differences in patient demographics or data collection practices. Similarly, in finance, the knowledge from credit scoring models in one region can be transferred to enhance models in another region.
@@ 2.3.3. Hierarchical Federated Learning
Hierarchical Federated Learning (HFL) (Abad et al., 2020; Feng et al., 2021; Liu et al., 2020; Briggs et al., 2020; Liu et al., 2022b) introduces a multi-tiered aggregation framework, significantly enhancing the scalability and efficiency of collaborative learning models. This system incorporates several levels of servers, from edge parameter servers to a final cloud parameter server, facilitating a more structured and efficient communication and aggregation pathway. By leveraging edge servers for initial aggregation (Liu et al., 2020), HFL mitigates the communication bottlenecks often encountered in traditional FL, enabling faster, more energy-efficient training processes. In response to the limitations observed in both FL and P2PL — where FL’s centralized model can pose privacy risks, and P2PL’s decentralized nature complicates integrity and consensus — HFL emerges as a powerful hybrid solution. It employs a hierarchical structure in which clients communicate updates to edge servers which then aggregate these updates before further communication among themselves or with a higher-level server. This architecture not only reduces the privacy concerns associated with a single central server but also addresses the communication and consensus challenges inherent in a fully decentralized setup.
2.3.4. Horizontal and Vertical Federated Learning
@@
2.3.5. Data Distribution Attributes and Differences
IID Data Distributions: In scenarios where data is considered IID, each client’s dataset is assumed to be drawn from the same probability distribution. This uniformity suggests that datasets across varied domains, such as those of a bank and a hospital, exhibit similar statistical properties and structures, despite their distinct contexts. Non-IID Data Distributions: On the other hand, non-IID data distributions reflect situations where data across clients originate from heterogeneous distributions. This diversity is rooted in demographic variances, regional characteristics, or the intrinsic nature of the collected data.
@@ 2.3.6. FedSGD and FedAvg: Core FL algorithms
The two standard FL algorithms are FedSGD and FedAvg (McMahan et al., 2017), mainly differ in the number of local iterations. FedSGD, the federated adaptation of stochastic gradient descent, operates on the principle of minimalism. It mandates that clients engage in a singular iteration of training on a localized data batch before transmitting the resultant gradient to a central server. This server, in turn, aggregates the gradients from all participating clients, adjusting the weights in proportion to each client’s contribution to the total dataset.
Advancing beyond FedSGD, FedAvg introduces multiple local iterations and/or epochs at each client before communicating updates to the server. These updates can comprise either the computed gradients over all local iterations or the updated model parameters themselves. The algorithm structure inherently minimizes the frequency of communication with the server, but also helps with privacy concerns by limiting the server exposure to initial and final model states only. This characteristic is surprisingly helpful in contexts where data sensitivity is paramount, as it effectively conceals the intermediate computational states from potential attackers.
2.4. Technical Challenges for Federated Learning
Federated learning expects computational and bandwidth resources to fully leverage the private data. Multiple federated learning algorithms have proposed solutions for efficient communication and/or compute resource scarcity. However, most of the deployed platforms are not built to address such requirements, making them the wrong choice for seamless, scalable federated learning. These platforms are often limited in resources, competing with other onboard functionalities, making scheduling an additional challenge. Earlier solutions were performing federated learning computations when the devices were idle, during the night and plugged in a power outlet. However, this can create biased models, where some of the data distributions are largely ignored.
2.4.1. Implications on Security of the Learning
Federated Learning introduces Security risks mainly for two reasons: the private data cannot be audited or curated, enabling “bad actors” to alter their data according to the end-goal, i.e. “training data attacks”. Also, the models are exposed during the communication process, introducing security risks while providing opportunities for backdoor attacks. These attacks aim at “influencing” the model parameters to a particular behavior, or simply degrading the overall model performance.
2.4.2. Data and Device Heterogeneity
Data and device heterogeneity can lead to biased models. The private data distributions and the available computational resources can vary widely between different clients, which leads to issues such as global model drift or the need to adopt multiple model architectures. FedAvg often shows performance disadvantages to highly mismatched data distributions, and several approaches have been proposed to address this issue. Besides any mismatch between distributions, the lack of labels during training can also produce suboptimal training conditions. This problem can be either mitigated by Knowledge Distillation (Cho et al., 2023), heuristics (Wang et al., 2021) or asking users to label their data (Augenstein et al., 2022). A basic assumption of the early FL algorithms was the local model structure is uniform across all clients. However, clients can differ in their computational resources and data distributions.
@@ 2.5. Threat model
For privacy attacks within the domain of FL, the the threat landscape is complex and varies significantly based on the nature of the attack. The architecture of FL inherently poses unique challenges in safeguarding privacy due to the decentralized nature of data processing.
Understanding the nuances of these threats is crucial for developing robust defense mechanisms. At the heart of FL, the central server embodies a critical point of vulnerability. It plays a dual role: distributing the learning model to clients and aggregating their updates. This centralized control positions the server as a potent target for attacks, making it the focal point for the strongest privacy threats in FL.
@@
Complexities of Misbehaving Clients: The threat landscape is further complicated by the variability in the behavior of misbehaving clients. Two critical factors need consideration: the proportion of clients engaged in malicious activities and the extent of their collaboration. Misbehaving clients may operate independently, presenting a scattered threat. However, the danger amplifies when they collude, orchestrating their attacks to maximize disruption or data extraction. This coordinated effort can significantly exacerbate the challenges in defending against privacy breaches and ensuring the integrity of the FL process (Sharma et al., 2023a).